# Ordinary primes in Hilbert modular varieties

### v1

@@ 2.3 Conjectures
And the analogue of the ‘ordinariness’ conjecture:
Conjecture 2.3.2
@@
We note that the right hand side of the conjectures is independent of $\lambda$ or $\ell$ by Corollary 2.2.4, and that the left hand side is independent of $\ell$ because the $IH^{n}(X,\mathbb{Q}_{\ell})$ form a strictly compatible system by Gabber [14] and Katz-Laumon [18, Th. 3.1.2]. Therefore Conjectures 2.3.1 and 2.3.2 are independent of the auxiliary prime $\ell$.
In case $X$ is also smooth, the conjectures are equivalent to those recalled in §1, since (1) the Newton polygons of the $\ell$-adic and crystalline Frobenius endomorphisms are the same by Katz and Messing [19], and (2) the Hodge-Tate polygon coincides with the Hodge polygon by Faltings [12].
@@ 3.1 Notation
We fix once and for all a rational prime $\ell$ that splits completely222we choose a split prime just for simplifying the exposition a little bit. The obvious analogues of Conjectures 2.3.1 and 2.3.2 are independent of $\ell$ for $M(f)$ defined below, since we still have a strictly compatible system of Galois representations. in $K_{f}$. For every nonarchimedean place $\lambda$ of $K_{f}$ dividing $\ell$, we denote by
@@
the associated semisimple, $K_{f}$-rational and integral Galois representation: See Deligne [10], Wiles [33], Taylor [29], Blasius and Rogawski [5] and the references therein. The Tate twist of its determinant $\det(\rho)(1)$ is a character of finite order.
@@ 3.2 Frobenius field and Ribet’s argument
Following Ribet [25], for every finite extension $F^{\prime}$ of $F$, we consider the Frobenius field:
@@
Then by Goursat’s Lemma (in the middle of the proof of Ribet [25, Th. 4.4.10]) the image of $\mathfrak{g}$ contains $\prod_{\lambda^{\circ}|\ell}\mathfrak{sl}_{2,\lambda^{\circ}}$, where $\lambda^{\circ}$ ranges over the primes of $K_{f}^{\circ}$ lying over $(\ell)$. Since the determinant on $G_{f,\lambda^{\circ}}^{\circ}$ is a dominant map onto $\mathbb{G}_{m}$, we get the desired equality. ∎
Definition 3.2.2
@@ 3.3 Variants of Sato-Tate equidistribution
We expect the strongest (SST) to be true; it fits into Serre’s general framework of Sato-Tate equidistribution [27, Chpt. 8], for almost all primes $\ell$ (depending on $f$).  [^4] The construction in §8.3 of op. cit., as stated, deals only with representations that come from the $\ell$-adic cohomology of algebraic varieties, but appears to use the condition only to the extent that they are rational and Hodge-Tate. In case $d$ is odd or the automorphic representation corresponding to $f$ has a discrete series at some finite prime, the Hodge-Tate condition (even the de Rham condition) for all $\ell$ follows from the motivic nature of the available constructions and theorems of Faltings [13], see Blasius and Rogawski [5]. In the general case, the Hodge-Tate condition is known for all but finitely many $\ell$: See Taylor [30], where they are shown to be (even) crystalline. Namely, the compact Lie group attached to $\rho_{f,\ell}|_{\widetilde{F^{\circ}}}$ by Serre is the product of $k_{f}^{\circ}$ replicas of $SU_{2}$ and the axioms $(A_{1})$ and $(A_{2})$ should hold.
When $t<k_{f}^{\circ}$, the conjunction of (RST) and ($t$-ST) does not imply (SST).
Remark 3.3.2
The condition (SST) is stronger than the (usual) Sato-Tate equidistribution theorems available at the moment (see [16], [3], [2]).
@@
The case of ($t$-ST), for $t\leq 2$, looks accessible, see Harris [15].
@@ 3.5 Interaction of $F$ and $K$: Slope and Bisection
$\sigma_{F}(K)=1-\frac{\lambda_{F}(K)}{[K:\mathbb{Q}]}\,\,\,\,\in[0,1]\cap\mathbb{Q}.$
@@ 3.6 Zariski density and Haar density (Serre)
This follows from Serre [27, Prop.5.12] applied to $Z:=\varphi^{-1}(S)$, which is a proper algebraic subset of $G$ by the assumptions and has Zariski density 00 by definition. ∎
@@ 4.1 Non CM case
The key fact (from Brylinski and Labesse [6]) that we use from the constructions (see Deligne [10], Wiles [33], Taylor [29] and Blasius and Rogawski [5] and the references therein) of the Galois representations associated with the $\{\sigma(f)\}$ is the following: the $\ell$-adic étale realisation $M(f)_{\ell}$ of $M(f)$ is the direct sum of the tensor inductions (see Curtis and Reiner [7, §80C]):
@@
This implies that there exist elements555transversals (coset representatives) of $\operatorname{Gal}(\overline{\mathbb{Q}}/\mathbb{Q})$ modulo $\operatorname{Gal}(\overline{\mathbb{Q}}/F)$ $g_{1},\cdots,g_{d}\in\operatorname{Gal}(\overline{\mathbb{Q}}/\mathbb{Q})$ such that, for any element $\gamma$ in the core [^6] Recall that the core of a subgroup $H$ in a group $G$ is the largest normal subgroup of $G$ contained in $H$. $\operatorname{Gal}(\overline{\mathbb{Q}}/\widetilde{F})$ of $\operatorname{Gal}(\overline{\mathbb{Q}}/F)$ in $\operatorname{Gal}(\overline{\mathbb{Q}}/\mathbb{Q})$, the action of $\gamma$ on $M(f)_{\ell}$ is given by
@@
and let $R_{\mathfrak{p}}=\{\alpha_{\mathfrak{p}},\beta_{\mathfrak{p}}\}\subset K^{\prime}$ be the roots. Each embedding $\sigma:K_{f}\longrightarrow\mathbb{Q}_{\ell}$ extends to (at most $2$) embeddings $\sigma^{\prime}:K^{\prime}\longrightarrow\overline{\mathbb{Q}}_{\ell}$, and the image $\sigma^{\prime}(R_{\mathfrak{p}})\subset\overline{\mathbb{Q}}_{\ell}$ is independent of the choice $\sigma^{\prime}$. Therefore we can unambiguously form the multiset of slopes777this may not have integral breaking points:
@@
In order to proceed further, we first recall (generalised Ramanujan-Petersson conjecture, see Taylor [31]):
Lemma 4.1.3
@@
Proof of Lemma. In case $d$ is odd or the automorphic representation $\pi_{f}$ corresponding to $f$ is a discrete series representation at some finite prime, this follows from the essentially motivic nature of some of the constructions, see Blasius and Rogawski [5], together with Deligne’s proof of the Weil conjectures [11].
This can be proved in the general case, and only with the a priori non motivic constructions of Wiles [33] and Taylor [29]. Note that by the description (4.1.1.1), the algebraic integers $\sigma^{\prime}(\alpha_{\mathfrak{p}}^{d})\,\mbox{ and }\,\sigma^{\prime}(\beta_{\mathfrak{p}}^{d})$ for any embedding $\sigma^{\prime}$ of $K^{\prime}$ into $\overline{\mathbb{Q}}_{\ell}$, are eigenvalues of $\operatorname{Frob}_{\mathfrak{p}}$ acting on the $IH^{d}$ of the Baily-Borel compactification $X^{BB}(\mathfrak{n})$ of the Hilbert modular variety.
Now this last variety admits a surjective, generically finite map from a projective smooth toroidal compactification over $\mathbb{Z}_{(p)}$ [24].
By the decomposition theorem for perverse sheaves [4], the $2$ algebraic integers therefore appear as eigenvalues of $\operatorname{Frob}_{\mathfrak{p}}$ in the $H^{d}$ of the projective smooth variety. Then by Deligne’s proof of the Weil conjectures, they have all the archimedean absolute values $=(\mathbb{N}\mathfrak{p})^{d/2}$. By taking the $d$-th root, we get the Lemma. $\square$
@@
where $\wedge^{2}(\iota)$ takes values in $GL_{6}$ and $\det$ takes values in $\mathbb{G}_{m}$.
@@
Now if the equlity holds in (4.1.3.5), then we necessarily have $e_{i}=1$ whenever $e_{i}>0$, and by Lemma 4.1.3 and the product formula
@@ 4.2 CM case
Since the restriction of $\rho_{f,\lambda}$ to $\widetilde{F^{\circ}}$ is then abelian and $K_{f}$-rational, by a theorem of Serre [26, Chpt. III, §3], augmented with a transcendence result of Waldschmidt [32] (see Henniart [17]), this restriction is locally algebraic (and semisimple by assumption). Then by a theorem of Ribet [25, §1.6] (which extends that of Serre [26, §III.2.3]), there exist (i) a $2$-dimensional $K_{f}$-rational vector subspace $V_{0}$ of $K_{f,\lambda}^{\oplus 2}$, (ii) a modulus $\mathfrak{m}$ of $\widetilde{F^{\circ}}$, and (iii) a rational representation
@@ 5.1 $F=\mathbb{Q}$
    If $K_{f}^{\circ}\neq K_{f}$, then part (2) of Theorem 4.1.1 provides a principally abundant set of ordinary primes. In case $K_{f}^{\circ}=K_{f}$, part (5) and the univariate (i.e. $t=1$) Sato-Tate equidistribution (proven in [16] and [3]) gives an abundant set of primes $p$ (of lower density $\geq 0.0794$) such that $k(p)\in\{0,1\}$; if in addition $f$ satisfies (RST), then part (6) will imply the abundance of ordinary ($k(p)=0$) primes.
@@ 5.2 $F=\mathbb{Q}(\sqrt{2})$
    For the $8$ forms $f$ in 329.2-c and 329.3-c, $[K_{f}:\mathbb{Q}]=4$, $\widetilde{K_{f}}/\mathbb{Q}$ has group $D_{8}$, and the image of $\operatorname{Gal}(\overline{\mathbb{Q}}/F)$ is a Klein subgroup just as in the exceptions (b) for $F=\mathbb{Q}(\sqrt{5})$, and $\sigma_{F}(K_{f})=1/2$.
@@ 6.1 Conjectures in a general setting
For the following statements, we assume that (a’) is true.
@@
Proposition 6.1.2
Assume that part (a) of Conjecture 6.1.1 is true and let $M$ be the André motive of $\mathfrak{M}$, and that the idempotent $e$ is an algebraic cycle, and let $\mathfrak{R}$ be the Grothendieck motive cut out by $e$ from $\mathfrak{M}$. Then:
@@
Then, on the one hand, by Katz and Messing [19, Th. 2], $\phi_{\mathrm{cris},\mathfrak{p}}^{[k(\mathfrak{p}):\mathbb{F}_{p}]}$ has the same (multiset of) eigenvalues as the $\ell$-adic Frobenius $\operatorname{Frob}_{\mathfrak{p}}$ on ${\mathfrak{R}}_{\ell}$. Therefore they have the same Newton polygons. This also proves (b).
@@
Now the statement (c) follows from Mazur’s theorem [20] applied to $\mathfrak{R}_{\mathrm{cris},\mathfrak{p}}$. In summary:
@@
For (2), let $\widetilde{F^{\prime}}$ be the normal closure of $F^{\prime}$ over $\mathbb{Q}$ and replace $F$ with $\widetilde{F^{\prime}}$, so that $\rho_{\ell}$ is abelian. Since $\rho_{\ell}$ is $\mathbb{Q}$-rational and Hodge-Tate, by a theorem of Serre [26, §III.2.3], it is associated with a $\mathbb{Q}$-rational representation $\phi_{0}:S_{\mathfrak{m}}\longrightarrow GL_{V_{0}}$, where $V_{0}$ is a $\mathbb{Q}$-form of $R_{\ell}$ and $\mathfrak{m}$ is a modulus of $F$.
@@
Since we already know part (b) of the conjecture, we can choose a rational prime $\ell$ that splits completely in $F$. Then we can identify the $\sigma$ with the embeddings of $F$ into $\mathbb{Q}_{\ell}$, once an embedding ${\iota}_{\ell}:\overline{\mathbb{Q}}\longrightarrow\overline{\mathbb{Q}}_{\ell}$ has been fixed. Under this identification, the multiset of the Hodge-Tate weights of $\rho_{\ell}$ at any $\ell$-adic place $\lambda:F\longrightarrow\overline{\mathbb{Q}}_{\ell}$ is ${\left\{n_{\sigma_{0,\lambda}(i)}\right\}_{i=1,\cdots,N}}$, where $\sigma_{0,\lambda}$ is the (unique) embedding such that $\iota_{\ell}\circ\sigma_{0,\lambda}=\lambda$, by [26, Prop. 2, §III.1.1].
On the other hand, let $\mathfrak{p}$ be any prime of $F$ lying over any rational prime $p\neq\ell$ that splits completely in $F$. We can also identify the embeddings $\sigma:F\longrightarrow\overline{\mathbb{Q}}$ with the embeddings into $\mathbb{Q}_{p}$, once we fix a $p$-adic place of $\iota_{p}:\overline{\mathbb{Q}}\longrightarrow\overline{\mathbb{Q}}_{p}$. Then the multiset of the $p$-adic valuations of $\rho_{\ell}(\operatorname{Frob}_{\mathfrak{p}})$ is given by ${\left\{n_{\sigma_{1,\mathfrak{p}}(i)}\right\}_{i=1,\cdots,N}}$, where $\iota_{p}\circ\sigma_{1,\mathfrak{p}}$ is the $p$-adic place $\mathfrak{p}$: See [26, Cor. 2., §II.3.4].

### v3

@@ 2.3 Conjectures
In case $X$ is also smooth, this is known to be true, by theorems of Katz and Messing [KM74] (comparing the $\ell$-adic Frobenius at $\mathfrak{p}$ with the crystalline Frobenius), of Mazur [Ma73] (showing that the Newton polygon lies on or above the Hodge polygon for the crystalline cohomology), and of Faltings [Fa88] (showing that the Hodge-Tate polygon is the same as the Hodge polygon).
And we formulate the analogue of the ‘ordinariness’ conjecture:
Conjecture 2.5.
@@
We note that the right hand side of the conjectures is independent of $\lambda$ or $\ell$ by Corollary 2.2.4, and that the left hand side is independent of $\ell$ because the $IH^{n}(X,\mathbb{Q}_{\ell})$ form a strictly compatible system by Gabber [Fu00] and Katz-Laumon [KL85, Th. 3.1.2]. Therefore Conjectures 2.4 and 2.5 are independent of the auxiliary prime $\ell$.
In case $X$ is also smooth, Conjecture 2.5 is equivalent to Conjecture 1.1 recalled in §1, since (1) the Newton polygons of the $\ell$-adic and crystalline Frobenius endomorphisms are the same by Katz and Messing [KM74], and (2) the Hodge-Tate polygon coincides with the Hodge polygon by Faltings [Fa88].
@@ 3.1 Notation
We note that the ordinariness in this context is equivalent to the following simple condition: $\mathfrak{p}$ is an ordinary prime (for $f$) if and only if $a_{\mathfrak{p}}$ is nonzero and does not belong to any prime ideal Weierstrass-$\wp$ of $\mathcal{O}_{K_{f}}$ lying over $(p)=\mathfrak{p}\cap\mathbb{Z}$.
We fix once and for all a rational prime $\ell$ that splits completely222we choose a split prime just for simplifying the exposition a little bit. The obvious analogues of Conjectures 2.4 and 2.5 are independent of $\ell$ for $M(f)$ defined below, since we still have a strictly compatible system of Galois representations. in $K_{f}$. For every nonarchimedean place $\lambda$ of $K_{f}$ dividing $\ell$, we denote by
@@
the associated semisimple, $K_{f}$-rational and integral Galois representation: See Deligne [D71], Ohta [Oh83], Carayol [C86], Wiles [Wi88], Taylor [T89], Blasius and Rogawski [BR93] and the references therein. The Tate twist of its determinant $\det(\rho)(1)$ is a character of finite order.
@@ 3.2 Frobenius field and Ribet’s argument
Following Ribet [Ri76], for every finite extension $F^{\prime}$ of $F$, we consider the Frobenius field:
@@
Then by Goursat’s Lemma (in the middle of the proof of Ribet [Ri76, Th. 4.4.10]) the image of $\mathfrak{g}$ contains $\prod_{\lambda^{\circ}|\ell}\mathfrak{sl}_{2,\lambda^{\circ}}$, where $\lambda^{\circ}$ ranges over the primes of $K_{f}^{\circ}$ lying over $(\ell)$. Since the determinant on $G_{f,\lambda^{\circ}}^{\circ}$ is a dominant map onto $\mathbb{\mathbf{G}}_{m}$, we get the desired equality. ∎
Definition 3.2.2.
@@ 3.3 Variants of Sato-Tate equidistribution
We expect the strongest (SST) to be true; it fits into Serre’s general framework of Sato-Tate equidistribution [Se12, Chpt. 8], for almost all primes $\ell$ (depending on $f$).  [^4] The construction in §8.3 of op. cit., as stated, deals only with representations that come from the $\ell$-adic cohomology of algebraic varieties, but appears to use the condition only to the extent that they are rational and Hodge-Tate. In case $d$ is odd or the automorphic representation corresponding to $f$ has a discrete series at some finite prime, the Hodge-Tate condition (even the de Rham condition) for all $\ell$ follows from the motivic nature of the available constructions and theorems of Faltings [Fa89], see Blasius and Rogawski [BR93]. In the general case, the Hodge-Tate condition is known for all but finitely many $\ell$: See Taylor [T95], where they are shown to be (even) crystalline. Namely, the compact Lie group attached to $\rho_{f,\ell}|_{\widetilde{F^{\circ}}}$ by Serre is the product of $k_{f}^{\circ}$ copies of $SU_{2}$ and the axioms $(A_{1})$ and $(A_{2})$ should hold.
When $t<k_{f}^{\circ}$, the mere conjunction of (RST) and ($t$-ST) does not imply (SST).
Remark 3.3.2.
The condition (SST) is stronger than the (usual) Sato-Tate equidistribution theorems available at the moment (see [HSBT10], [BLGHT11], [BLGG11]).
@@
The case of ($t$-ST), for $t\leq 2$, looks accessible, see Harris [Ha09].
@@ 3.5 Interaction of $F$ and $K$: Slope and Bisection
$\sigma_{F}(K)=1-\frac{\lambda_{F}(K)}{[K:\mathbb{Q}]}\,\,\,\,\in[0,1)\cap\mathbb{Q}.$
@@ 3.6 Zariski density and Haar density (Serre)
This follows from Serre [Se12, Prop.5.12] applied to $Z:=\varphi^{-1}(S)$, which is a proper algebraic subset of $G$ by the assumptions and has Zariski density 00 by definition. ∎
@@ 4.1 Non CM case
Denote by $M(f)$ the André motive (see Proposition 2.2.6), whose realisations give the part of the intersection cohomology of the Hilbert modular variety corresponding to $\{\sigma(f)\}_{\sigma}$, where $\sigma$ ranges over all the embeddings of $K_{f}$ into $\overline{\mathbb{Q}}$.
@@
The key fact (from Brylinski and Labesse [BL84]) that we use from the constructions (see Deligne [D71], Ohta [Oh83], Carayol [C86], Wiles [Wi88], Taylor [T89] and Blasius and Rogawski [BR93] and the references therein) of the Galois representations associated with the $\{\sigma(f)\}$ is the following: the $\ell$-adic étale realisation $M(f)_{\ell}$ of $M(f)$ is the direct sum of the tensor inductions (see Curtis and Reiner [CR, §80C]):
@@
This implies that for transversals (coset representatives) $g_{1},\cdots,g_{d}\in\operatorname{Gal}(\overline{\mathbb{Q}}/\mathbb{Q})$ of $\operatorname{Gal}(\overline{\mathbb{Q}}/\mathbb{Q})$ modulo $\operatorname{Gal}(\overline{\mathbb{Q}}/F)$, we have: For any element $\gamma$ in the finite index normal subgroup $\operatorname{Gal}(\overline{\mathbb{Q}}/\widetilde{F})$ of $\operatorname{Gal}(\overline{\mathbb{Q}}/F)$, the action of $\gamma$ on $M(f)_{\ell}$ is given by
@@
and let $R_{\mathfrak{p}}=\{\alpha_{\mathfrak{p}},\beta_{\mathfrak{p}}\}\subset K^{\prime}$ be the roots. Each embedding $\sigma:K_{f}\longrightarrow\mathbb{Q}_{\ell}$ extends to (at most $2$) embeddings $\sigma^{\prime}:K^{\prime}\longrightarrow\overline{\mathbb{Q}}_{\ell}$, and the image $\sigma^{\prime}(R_{\mathfrak{p}})\subset\overline{\mathbb{Q}}_{\ell}$ is independent of the choice $\sigma^{\prime}$. Therefore we can unambiguously form the multiset of slopes555this may not have integral breaking points:
@@
In order to proceed further, we first recall the following known fact (generalised Ramanujan-Petersson conjecture, see Taylor [T95] and Blasius[B06]):
Lemma 4.1.3.
@@
Proof of Lemma. In case $d$ is odd or the automorphic representation $\pi_{f}$ corresponding to $f$ is a discrete series representation at some finite prime, this follows from the essentially motivic nature of some of the constructions, see Blasius and Rogawski [BR93], together with Deligne’s proof of the Weil conjectures [D74].
This can be proved in the general case, and only with the a priori non motivic constructions of Wiles [Wi88] and Taylor [T89]. Note that by the description (4.1.1.1), the algebraic integers $\sigma^{\prime}(\alpha_{\mathfrak{p}}^{d})\,\mbox{ and }\,\sigma^{\prime}(\beta_{\mathfrak{p}}^{d})$ for any embedding $\sigma^{\prime}$ of $K^{\prime}$ into $\overline{\mathbb{Q}}_{\ell}$, are eigenvalues of $\operatorname{Frob}_{\mathfrak{p}}$ acting on the $IH^{d}$ of the Baily-Borel compactification $X^{BB}(\mathfrak{n})$ of the Hilbert modular variety.
Now this last variety admits a surjective, generically finite map from a projective smooth toroidal compactification over $\mathbb{Z}_{(p)}$ [Ra78].
By the decomposition theorem for perverse sheaves [BBD], the $2$ algebraic integers therefore appear as eigenvalues of $\operatorname{Frob}_{\mathfrak{p}}$ in the $H^{d}$ of the projective smooth variety. Then by Deligne’s proof of the Weil conjectures, they have all the archimedean absolute values $=(\mathbb{N}\mathfrak{p})^{d/2}$. By taking the $d$-th root, we get the Lemma. $\square$
@@
where $\wedge^{2}(\iota)$ takes values in $GL_{6}$ and $\det$ takes values in $\mathbb{\mathbf{G}}_{m}$.
@@
Now if the equality holds in (4.1.3.5), then we necessarily have $e_{i}=1$ whenever $e_{i}>0$, and by Lemma 4.1.3 and the product formula
@@ 4.2 CM case
Since the restriction of $\rho_{f,\lambda}$ to $\widetilde{F^{\circ}}$ is then abelian and $K_{f}$-rational, by a theorem of Serre [Se98, Chpt. III, §3], augmented with a transcendence result of Waldschmidt [Ws81] (see Henniart [He82]), this restriction is locally algebraic (and semisimple by assumption). Then by a theorem of Ribet [Ri76, §1.6] (which extends that of Serre [Se98, §III.2.3]), there exist (i) a $2$-dimensional $K_{f}$-rational vector subspace $V_{0}$ of $K_{f,\lambda}^{\oplus 2}$, (ii) a modulus $\mathfrak{m}$ of $\widetilde{F^{\circ}}$, and (iii) a rational representation
@@ 5.1 $F=\mathbb{Q}$
    If $K_{f}^{\circ}\neq K_{f}$, then part (2) of Theorem 4.1.1 provides a principally abundant set of ordinary primes. In case $K_{f}^{\circ}=K_{f}$, part (5) and the univariate (i.e. $t=1$) Sato-Tate equidistribution (proven in [HSBT10] and [BLGHT11]) gives an abundant set of primes $p$ (of lower density $\geq 0.0794$) such that $k(p)\in\{0,1\}$; if in addition $f$ satisfies (RST), then part (6) will imply the abundance of ordinary ($k(p)=0$) primes.
@@ 5.2 $F=\mathbb{Q}(\sqrt{2})$
    For the $8$ forms $f$ in 329.2-c and 329.3-c, $[K_{f}:\mathbb{Q}]=4$, $\widetilde{K_{f}}/\mathbb{Q}$ has group $D_{8}$, the image of $\operatorname{Gal}(\overline{\mathbb{Q}}/F)$ is a Klein subgroup, and $\sigma_{F}(K_{f})=1/2$.
@@ 6.1 Conjectures in a general setting
For the following statements, we assume that (a’) is true. [^7] It appears that (a’) follows from the construction of pure Nori motives realising the $\ell$-adic intersection cohomology groups, due to Ivorra and Morel [IM19, §6].
@@
Proposition 6.2.1.
Assume that part (a) of Conjecture 6.2 is true and let $M$ be the André motive of $\mathfrak{M}$, and that the idempotent $e$ is an algebraic cycle, and let $\mathfrak{R}$ be the Grothendieck motive cut out by $e$ from $\mathfrak{M}$. Then:
@@
Then, on the one hand, by Katz and Messing [KM74, Th. 2], $\phi_{\mathrm{cris},\mathfrak{p}}^{[k(\mathfrak{p}):\mathbb{F}_{p}]}$ has the same (multiset of) eigenvalues as the $\ell$-adic Frobenius $\operatorname{Frob}_{\mathfrak{p}}$ on ${\mathfrak{R}}_{\ell}$. Therefore they have the same Newton polygons. This also proves (b).
@@
Now the statement (c) follows from Mazur’s theorem [Ma73] applied to $\mathfrak{R}_{\mathrm{cris},\mathfrak{p}}$. In summary:
@@
For (2), let $\widetilde{F^{\prime}}$ be the normal closure of $F^{\prime}$ over $\mathbb{Q}$ and replace $F$ with $\widetilde{F^{\prime}}$, so that $\rho_{\ell}$ is abelian. Since $\rho_{\ell}$ is $\mathbb{Q}$-rational and Hodge-Tate, by a theorem of Serre [Se98, §III.2.3], it is associated with a $\mathbb{Q}$-rational representation $\phi_{0}:S_{\mathfrak{m}}\longrightarrow GL_{V_{0}}$, where $V_{0}$ is a $\mathbb{Q}$-form of $R_{\ell}$ and $\mathfrak{m}$ is a modulus of $F$.
@@
Since we already know part (b) of the conjecture, we can choose a rational prime $\ell$ that splits completely in $F$. Then we can identify the $\sigma$ with the embeddings of $F$ into $\mathbb{Q}_{\ell}$, once an embedding ${\iota}_{\ell}:\overline{\mathbb{Q}}\longrightarrow\overline{\mathbb{Q}}_{\ell}$ has been fixed. Under this identification, the multiset of the Hodge-Tate weights of $\rho_{\ell}$ at any $\ell$-adic place $\lambda:F\longrightarrow\overline{\mathbb{Q}}_{\ell}$ is ${\left\{n_{\sigma_{0,\lambda}(i)}\right\}_{i=1,\cdots,N}}$, where $\sigma_{0,\lambda}$ is the (unique) embedding such that $\iota_{\ell}\circ\sigma_{0,\lambda}=\lambda$, by [Se98, Prop. 2, §III.1.1].
On the other hand, let $\mathfrak{p}$ be any prime of $F$ lying over any rational prime $p\neq\ell$ that splits completely in $F$. We can also identify the embeddings $\sigma:F\longrightarrow\overline{\mathbb{Q}}$ with the embeddings into $\mathbb{Q}_{p}$, once we fix a $p$-adic place of $\iota_{p}:\overline{\mathbb{Q}}\longrightarrow\overline{\mathbb{Q}}_{p}$. Then the multiset of the $p$-adic valuations of $\rho_{\ell}(\operatorname{Frob}_{\mathfrak{p}})$ is given by ${\left\{n_{\sigma_{1,\mathfrak{p}}(i)}\right\}_{i=1,\cdots,N}}$, where $\iota_{p}\circ\sigma_{1,\mathfrak{p}}$ is the $p$-adic place $\mathfrak{p}$: See [Se98, Cor. 2., §II.3.4].